OpenStudy (anonymous):

sin squared theta - 5 sin theta - 6 = 0

OpenStudy (anonymous):

I was wondering how to solve this problem?

OpenStudy (anonymous):

solve \[x^2-5x-6=0\] for \(x\) and then replace \(x\) by \(\sin(\theta)\)

OpenStudy (anonymous):

(Enter your answers as a comma-separated list. Let k be any integer. Round terms to three decimal places where appropriate. If there is no solution, enter NO SOLUTION.)

OpenStudy (anonymous):

Is the quadratic formula the best way to find the answer?